OpenStudy (anonymous):

Is anyone here who is expert in relations and functions and inverse trigonometry ? Please I want to ask some questions and have to submit their solutions tomorrow. Please answer.

OpenStudy (anonymous):

1st question is : Let A = N X N and let * be a binary operation on A defined by (a,b) * (c,d) = (ac, bd). Show that (i) (A,*) is commutative (ii) (A,*) is associative Find the identity element , if any, in A.

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

ok commutative is?

OpenStudy (anonymous):

if changing the order of the operands does not change the end result

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

f(A,*)=f(*,A) that's what we have to proof

OpenStudy (anonymous):

yes

OpenStudy (zarkon):

(c,d)*(a,b)=?

OpenStudy (anonymous):

(ca, db) = (ac, bd) = (a,b) * (c,d)

OpenStudy (nilankshi):

good kush

OpenStudy (anonymous):

This is what you want to say @ Zarkon Right ?

OpenStudy (zarkon):

yes...that is all you need to show for (i)

OpenStudy (anonymous):

sorry comp froze, have an ancient thing on mw yes changing order didn't change the end result right

OpenStudy (anonymous):

in ii part

OpenStudy (zarkon):

show... [(a,b) * (c,d) ]*(e,f)=(a,b) *[ (c,d) *(e,f)]

OpenStudy (anonymous):

does it matter which u multiply first is the question u should be asking

OpenStudy (anonymous):

is this queston for me

OpenStudy (anonymous):

oh like zarkon said

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wait zarkon do we introduce new variables?

OpenStudy (zarkon):

you just pick any three elements from the set NXN

OpenStudy (anonymous):

[(a,b) * (c,d)]*(e,f) = [(ac,bd)]*(e,f) =(ace,bdf) =(a,b)*(ce,df) =(a,b)*[(c,e)*(d,f)] Proved

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Is this what we have to show ?

OpenStudy (anonymous):

i think so

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

Find the identity element , if any, in A.

OpenStudy (zarkon):

is N the set of natural numbers?

OpenStudy (anonymous):

WILL THE ANSWER FOR THIS BE (1,1)

OpenStudy (anonymous):

we say any another element given relation, prove that (a,b)*(c,d)*(e,f) where order of multiplication doesnt matter and ef is another point in the set

OpenStudy (zarkon):

if N is the natural numbers then yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

lol this isnt hs math

OpenStudy (anonymous):

is it?

OpenStudy (zarkon):

sure it is

OpenStudy (anonymous):

can u help me in some more questions

OpenStudy (anonymous):

really zarkon

OpenStudy (zarkon):

basic stuff you would see in a HS discrete math class. or a general humanities math class in college

OpenStudy (anonymous):

sure post them again though

OpenStudy (anonymous):

hmm they dont have tht clas in my school

OpenStudy (anonymous):

okay wait i will post them

OpenStudy (nilankshi):

hahd is a good teacher

OpenStudy (anonymous):

i learn out of interest @nil loooooool prob not this was all zarkon

OpenStudy (nilankshi):

yes

OpenStudy (nilankshi):

hahd u are in which class

OpenStudy (anonymous):

gr.11

OpenStudy (anonymous):

hw abt u zarkon?

OpenStudy (zarkon):

what?

OpenStudy (anonymous):

gr.12?

OpenStudy (anonymous):

what grade are u in?

OpenStudy (nilankshi):

u ar from

OpenStudy (anonymous):

Canada

OpenStudy (zarkon):

lol...I'm a Mathematics Professor

OpenStudy (anonymous):

WOOOOOOW, k respect right there

OpenStudy (nilankshi):

PROFESSOR

OpenStudy (anonymous):

zarkon where do u teach?

OpenStudy (zarkon):

In Pennsylvania

OpenStudy (nilankshi):

Then all my probs by u

OpenStudy (anonymous):

yh american math curriculum is better than canada's

OpenStudy (anonymous):

http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e509b4e0b8b958804ac0716 go to this link to answer the question